Explain and define the Earnings Per Share, and using the following data, calculate the EPS for Rotor and Mora Valve Manufacturing Companies: Amounts in US$ 000’s Rotor Mora Net Profit After Tax 50,000 40,000 If both companies share the same Shares Outstanding of 100,000 in the First Half of the year, then Rotor Reported 200,000 Shares at Year End, While Mora Reported 300,000 Shares at the Same Period.
Answer #1

Earning per share is a financial ratio which shows the profitability to share holders. It is a ratio of earning available to share holders and weighted average of outstanding shares. It can be computed with following formula:

Net Income EPS = weighted average of outstanding share

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

Net Profit after tax Weighted average of outstanding shares Rotor $50,000,000.00 150,000 $333.33 Mora $40,000,000.00 200,000

Cell reference -

В C D E Rotor Net Profit after tax 50000000 Weighted average of outstanding shares =E12 EPS =C3/04 Mora 40000000 =E19 =D3/D4

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.